OpenStudy (anonymous):

For 2x – 3y = –10, when x = 4, What is the value of y? Please help

OpenStudy (anonymous):

plug in the value of 4 into x, and what do you get?

OpenStudy (anonymous):

5?

OpenStudy (anonymous):

try one more time, 2(4)-3y= -10

OpenStudy (anonymous):

i got nothing... i'm really bad at solving equations :/

OpenStudy (anonymous):

no worries, if you plug 4 into x, you're basically saying 2 times 4 as is shown by 2(4) - 3y = -10, or 8-3y=-10 do you see how i got that?

OpenStudy (anonymous):

ya i got that part but now idk what to do after

OpenStudy (anonymous):

next you want to isolate the variables.. basically get the y by itself on one side of the equation. so, you want to move the 8 over to the other side. how would you dod that?

OpenStudy (anonymous):

do that*

OpenStudy (anonymous):

move the 8 to the other side so it would be 3y=-10-8?

OpenStudy (anonymous):

exactly! or 3y= - 18, now you want to get y by itself, do you know how you would do that?

OpenStudy (anonymous):

divde 3/y=-3/18

OpenStudy (anonymous):

-6!

OpenStudy (anonymous):

exactly!!! just keep in mind though, that you're putting the denominator on the bottom, like so: y/3 = -18/3 cuz in the future it might get confusing ^.^, but good job!

OpenStudy (anonymous):

oh wait, its divide by -3, so it would actually be 6, sorry!

OpenStudy (anonymous):

y/-3 = -18/-3 = 6

OpenStudy (anonymous):

oooo ok

OpenStudy (anonymous):

do you know what type of line is the graph of y=7?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

does that process make sense? when you solve for a variable, you're just trying to get it alone. so you're trying to graph y=7? any time you're graphing one number per variable, it's just a straight line crossing the axis at that variable, so for example to graph x =5 it would be  if you're trying to graph y = 5, it would be :